# Kalb-Ramond field, black holes and black strings in (2 + 1)D

### v4

@@ Appendix A. Invariant metric on $AdS_{3}$
We simply replace $\theta$ with $\overline{\theta}\in{\tenrm I\kern-1.79993ptR}$.
@@ Appendix B. Bosonic T-duality
We briefly review T-duality in Type II bosonic string theory. See [[38]38[39]--[40]40,,  [41]41] for detailed discussions. See also Appendix A in [1].
@@
where $G_{mn}=G_{nm}$ are the components of the metric $G$ and $B_{mn}=-B_{mn}$ are the components of the Kalb-Ramond two-form $B$. $z=\tau+\sigma$ and ${\overline{z}}=\tau-\sigma$ are the world-sheet coordinates. We omitted a term linear in the dilaton field. The dilaton $\Phi$ enters at the quantum level [[42]42,,[43]43].
@@
is a symmetry of the action, where $c$ is a constant. Therefore, we can couple the theory to a gauge field $A_{a}$ on the worldsheet. The gauged Lagrangian density is
@@
We assume $X^{1}$ is periodic and compact, i.e., at fixed worldsheet time it parametrizes a circle. Thus, the symmetry (B.2) is a global $U(1)$ (zero-form) symmetry. We fix the gauge by assuming $X^{1}={\tenrm const.}$. This gives
@@
The relations (B.8), (B.10) and (B.11) are also known as the Buscher rules. From the worldsheet perspective, T-dulity is an exact equivalence of CFTs.
@@ Appendix C. The Brown-York quasi-local energy and mass
The Brown-York definition of energy and mass is based on the Hamilton-Jacobi formulation of general relativity. It requires foliating the spacetime into a one-parameter family of space-like hyper-surfaces and thus defining a time direction. It also requires adding the Gibbons-Hawking-York (GHY) boundary term to the usual action for the gravitational field [[44]44,,[45]45]. The boundary term ensures that only the metric is fixed at the boundary. It avoids the need to fix the first derivatives of the metric.
@@
where $A,B,C$ and $Z$ are functions of only the radial coordinate $r$. The coordinate $\theta$ is an angle variable and it has $2\pi$ periodicity. The spacetime contains no inner boundaries. It can have event horizons [4].22 In the case the spacetime contains other null boundaries we include appropriate terms in the action functional that contribute on these boundaries. This will not concern us in this paper. We refer to [46] and references therein for such cases. We assume the metric is minimally coupled to a scalar. Therefore, the scalar action contains no derivatives of the metric. We also assume the scalar depends only on $r$.
Let $u^{\mu}$ denote the future pointing time-like unit normal to the $t={\tenrm constant}$ two-space denoted by $\Sigma_{t}$. $t$ labels the two-spaces $\Sigma_{t}$ and it defines the time direction or coordinate. This particular slicing is the usual ADM slicing of spacetime [47]. We denote the (intrinsic) metric on $\Sigma_{t}$ by $h_{ij}$.23 For convenience, here and below we use lower latin letters $i=t,r,\theta$ to denote the intrinsic metrics and we use greek letters $\mu=t,r,\theta$ to denote the induced metrics. We denote the boundary of $\Sigma_{t}$ by $B_{t}$. The world-line or history of $B_{t}$ is denoted ${}^{2}B$. Let $n^{\mu}$ is the outward pointing space-like unit normal to the two-boundary ${}^{2}B$. For simplicity, we assume on the time-like boundary ${}^{2}B$ that $n^{\mu}u_{\mu}|_{{}^{2}B}=0$. That is, the surface $\Sigma_{t}$ meets the boundary ${}^{2}B$ orthogonally. This is not however essential or necessary for the analysis below [[48]48]. This is in part because we are ultimately interested in the case ${}^{2}B$ is at radial infinity where the spacetime (C.1) is (conformally) flat. We denote the metric on ${}^{2}B$ by $\gamma_{ij}$. We denote the metric on $B_{t}$ which is the intersection of ${}^{2}B$ and $\Sigma_{t}$ by $\sigma_{ij}$.
@@
We now give the conserved charges in terms of the components of the metric. The energy associated with the Killing field $\xi^{i}=u^{i}=\gamma_{i}^{\mu}u_{\mu}$24 Note ${\cal D}_{(i}\xi_{j)}=\gamma_{(i}^{\lambda}\nabla_{\lambda}\gamma_{j)}^{\mu}u_{\mu}=0$. is given by
@@
In gravity theories physical quantities such as energy are defined only with respect to a reference. Thus, we must specify a reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$, where $g_{0}$ denotes the metric and $\phi_{0}$ collectively denotes the matter fields on ${\cal M}_{0}$. $\Lambda$ is possible cosmological constant.
@@
vanish smoothly at the boundary $B_{t}$ [3]. Note $B_{t}$ is the boundary of the $t={\tenrm constant}$ spacetime slice. In the case ${\cal M}(g,\phi;\Lambda)$ is asymptotically flat spacetime, the reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$ is usually taken to be the Minkowski spacetime. In this paper in all the cases we study the reference metrics are conformally flat spacetimes.
@@
The reference metrics $g_{0}$ that we consider in this paper have $Z_{0}={\tenrm constant}$. Accordingly, since $\partial_{r}Z_{0}=0$, the (quasi-local) angular momentum remains unchanged. It is still given by (C.21). The (quasi-local) energy is given by26 In general it is quasi-local in the sense that it is local only on the (hyper-)surface $\Sigma_{t}$, i.e., the $t={\tenrm constant}$ spacetime slice. It follows, in $(2+1)$ dimensions, it is defined as a local quantity.
@@ Appendix D. Kruskal coordinates
The metrics which we study in this paper are in general of the form $ds^{2}=\Delta^{2}d{\widehat{s}}^{2}$, where the factor $\Delta^{2}$ is (manifestly) positive and regular everywhere.27 Except at a curvature singularity. The line element $d{\widehat{s}}^{2}$ is of two types. One type has the form (see, e.g., (3.15) and (4.14))
@@
and the other has (in general with different $f,g$ and $\Delta$) the form (see, e.g., (3.24) and (3.48))
@@
The functions $f,g$ and $\Delta$ depend only on $\rho$. $\phi$ is angular coordinate. The function $f$ has either one or two simple zeros or poles. The scalar curvature is finite at the zero(s). At large $\rho$ both are conformally flat. Thus, each of their Penrose diagrams is similar at large $\rho$ to that of a flat spacetime.
@@
In the large $x$ limit we get $x^{\star}=2|c_{2}|c_{1}^{2}\ln x+h(c_{1},c_{2},\lambda)$.28 The BTZ solution is obtained by setting $\mu=1$, i.e., $c_{2}=0$. For $c_{2}=0$, in the large $x$ limit, we instead obatin $x^{\star}=-c_{1}^{2}/x+{\cal O}(x^{-3})$. The function $h$ in particular satisfies $h(c_{1},0,\lambda)=0$. In the large $x$ limit and at constant $\varphi$ the metric (D.3) up to an overall positive factor equals $du^{2}-dv^{2}+[\lambda^{2}/(u^{2}-v^{2})](udv-vdu)^{2}$. Therefore, outgoing and ingoing radial null lines end on future and past null infinities, respectively.

### v5

@@ Appendix A. Invariant metric on $AdS_{3}$
We simple replace $\theta$ with ${\overline{\theta}}\in{\tenrm I\kern-1.79993ptR}$.
@@ Appendix B. Bosonic T-duality
We briefly review T-duality in Type II bosonic string theory. See [[41]41[42]--[43]43,,  [44]44] for detailed discussions. See also Appendix A in [1].
@@
where $G_{mn}=G_{nm}$ are the components of the metric $G$ and $B_{mn}=-B_{mn}$ are the components of the Kalb-Ramond two-form $B$. $z=\tau+\sigma$ and ${\overline{z}}=\tau-\sigma$ are the world-sheet coordinates. We omitted a term linear in the dilaton field. The dilaton $\Phi$ enters at the quantum level [[45]45,,[46]46].
@@
is a symmetry of the action, where $c$ is a constant. Therefore, we can couple the theory to a gauge field $A_{a}$ on the world-sheet. The gauged Lagrangian density is
@@
We assume $X^{1}$ is periodic and compact, i.e., at fixed world-sheet time it parametrizes a circle. Thus, the symmetry (B.2) is a global $U(1)$ (zero-form) symmetry. We fix the gauge by assuming $X^{1}={\tenrm const.}$. This gives
@@
The relations (B.8), (B.10) and (B.11) are also known as the Buscher rules. From the world-sheet perspective, T-duality is an exact equivalence of CFTs.
@@ Appendix C. The Brown-York quasi-local energy and mass
The Brown-York definition of energy and mass is based on the Hamilton-Jacobi formulation of general relativity. It requires foliating the spacetime into a one-parameter family of space-like hyper-surfaces and thus requires defining a time direction. It also requires adding the Gibbons-Hawking-York (GHY) boundary term to the usual action for the gravitational field [[47]47,,[48]48]. The boundary term ensures that only the metric is fixed at the boundary. It avoids the need to fix the first derivatives of the metric.
@@
where $A,B,C$ and $Z$ are functions of only the radial coordinate $r$. The coordinate $\theta$ is an angle variable and it has $2\pi$ periodicity. The spacetime contains no inner boundaries. It can have event horizons [4].30 In the case the spacetime contains other null boundaries we include appropriate terms in the action functional that contribute on these boundaries. This will not concern us in this paper. We refer to [49] and references therein for such cases. We assume the metric is minimally coupled to a scalar. Therefore, the scalar action contains no derivatives of the metric. We also assume the scalar depends only on $r$.
Let $u^{\mu}$ denote the future pointing time-like unit normal to the $t={\tenrm constant}$ two-space denoted by $\Sigma_{t}$. $t$ labels the two-spaces $\Sigma_{t}$ and it defines the time direction or coordinate. This particular slicing is the usual ADM slicing of spacetime [50]. We denote the (intrinsic) metric on $\Sigma_{t}$ by $h_{ij}$.31 For convenience, here and below we use lower latin letters $i=t,r,\theta$ to denote the intrinsic metrics and we use greek letters $\mu=t,r,\theta$ to denote the induced metrics. We denote the boundary of $\Sigma_{t}$ by $B_{t}$. The world-line or history of $B_{t}$ is denoted ${}^{2}B$. Let $n^{\mu}$ is the outward pointing space-like unit normal to the two-boundary ${}^{2}B$. For simplicity, we assume on the time-like boundary ${}^{2}B$ that $n^{\mu}u_{\mu}|_{{}^{2}B}=0$. That is, the surface $\Sigma_{t}$ meets the boundary ${}^{2}B$ orthogonally. This is not however essential or necessary for the analysis below [[51]51]. This is in part because we are ultimately interested in the case ${}^{2}B$ is at radial infinity where the spacetime (C.1) is (conformally) flat. We denote the metric on ${}^{2}B$ by $\gamma_{ij}$. We denote the metric on $B_{t}$ which is the intersection of ${}^{2}B$ and $\Sigma_{t}$ by $\sigma_{ij}$.
@@
We now give the conserved charges in terms of the components of the metric. The energy associated with the Killing field $\xi^{i}=u^{i}=\gamma_{i}^{\mu}u_{\mu}$32 Note ${\cal D}_{(i}\xi_{j)}=\gamma_{(i}^{\lambda}\nabla_{\lambda}\gamma_{j)}^{\mu}u_{\mu}=0$. is given by
@@
In gravity theories physical quantities such as energy are defined only with respect to a reference. Thus, we must specify a reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$, where $g_{0}$ denotes the metric and $\phi_{0}$ collectively denotes the matter fields on ${\cal M}_{0}$. $\Lambda$ is a possible cosmological constant.
@@
vanish smoothly at the boundary $B_{t}$ [3]. Note $B_{t}$ is the boundary of the $t={\tenrm constant}$ spacetime slice. In the case ${\cal M}(g,\phi;\Lambda)$ is asymptotically flat spacetime, the reference spacetime ${\cal M}_{0}(g_{0},\phi_{0};\Lambda)$ is usually taken to be the Minkowski spacetime. In the paper in all the cases we study the reference metrics are conformally flat spacetimes.
@@
The reference metrics $g_{0}$ that we consider in this paper have $Z_{0}={\tenrm constant}$. Accordingly, since $\partial_{r}Z_{0}=0$, the (quasi-local) angular momentum remains unchanged. It is still given by (C.21). The (quasi-local) energy is given by34 In general it is quasi-local in the sense that it is local only on the (hyper-)surface $\Sigma_{t}$, i.e., the $t={\tenrm constant}$ spacetime slice. It follows, in $(2+1)$ dimensions, it is defined as a local quantity.
@@ Appendix D. Kruskal coordinates
The metrics which we study in this paper are in general of the form $ds^{2}=\Delta^{2}d{\widehat{s}}^{2}$, where the factor $\Delta^{2}$ is (manifestly) positive and regular everywhere.35 Except at a curvature singularity. The line element $d{\widehat{s}}^{2}$ is of two types. One type has the form (see, e.g., (3.15) (with (3.27)) and (4.13))
@@
and the other has (in general with different $f,g$ and $\Delta$) the form (see, e.g., (3.20), (3.31) and (3.56))
@@
The functions $f,g$ and $\Delta$ depend only on $\rho$. $\phi$ is an angular coordinate. The function $f$ has either one or two simple zeros or poles. The scalar curvature is finite at the zero(s). At large $\rho$ both are conformally flat. Thus, each of their Penrose diagrams is similar at large $\rho$ to that of a flat spacetime.
@@
In the large $x$ limit we get $x^{\star}=2|c_{2}|c_{1}^{2}\ln x+h(c_{1},c_{2},\lambda)$.36 The BTZ solution is obtained by setting $\mu=1$, i.e., $c_{2}=0$. For $c_{2}=0$, in the large $x$ limit, we instead obtain $x^{\star}=-c_{1}^{2}/x+{\cal O}(x^{-3})$. The function $h$ in particular satisfies $h(c_{1},0,\lambda)=0$. In the large $x$ limit and at constant $\varphi$ the metric (D.3) up to an overall positive factor equals $du^{2}-dv^{2}+[\lambda^{2}/(u^{2}-v^{2})](udv-vdu)^{2}$. Therefore, outgoing and ingoing radial null lines end on future and past null infinities, respectively.